OpenStudy (anonymous):

what is the equation in standard form of the line which passes through (-2,6) and has a slope of -1

OpenStudy (anonymous):

y=mx+b 6=-1(-2)+b 6=2+b b=4 So, equation of a line is: y=-1x+4??? Is this what you need?

OpenStudy (anonymous):

x-y=-4 x-y=4 x+y=4 x+y=-4

OpenStudy (anonymous):

Standard form of a line is Ax + Bx = C x+y=4